(4.1) Consider the point A = (−1,0,1), B = (0,−2,3), and C = (−4,4,1) to be vertices of a triangle ∆. Evaluate all side lengths of ∆. Let ∆ be the triangle with vertices the points P = (3,1,−1), Q = (2,0,3) and R = (1,1,1). (4.2)Determine whether ∆ is a right angle triangle. If it is not, explain with reason, why?

(4.3)Let u =< 0,1,1 >,v =< 2,2,0 >and w =< −1,1,0 >bethreevectorsin standard form.

(a) Determine which two vectors form a right angle triangle?

(b)Find θ := uw, the angel between the given two vectors.

(4.4) Let x < 0. Find the vector n =< x,y,z > that is orthogonal to all three vectors u =<1,1,−2 >,v =< −1,2,0 >and w =< −1,0,1 >.

(4.5) Find a unit vector that is orthogonal to both u =< 0,−1,−1 > and v =< 1,0,−1 >


1
Expert's answer
2021-06-30T09:22:23-0400

QUESTION 4.1

As we know, the distance between points "A\left(x_A,y_A,z_A\right)" and "B\left(x_B,y_B,z_B\right)" given by their Cartesian coordinates can be calculated by the formula



"AB=\sqrt{\left(x_A-x_B\right)^2+\left(y_A-y_B\right)^2+\left(z_A-z_B\right)^2}"

In our case,


"\left\{\begin{array}{l}
A\left(-1,0,1\right)\\
B\left(0,-2,3\right)\\
C\left(-4,4,1\right)\\
\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
AB=\sqrt{\left(0+1\right)^2+\left(0+2\right)^2+\left(3-1\right)^2}\\[0.3cm]
BC=\sqrt{\left(0+4\right)^2+\left(4+2\right)^2+\left(3-1\right)^2}\\[0.3cm]
AC=\sqrt{\left(-1+4\right)^2+\left(4-0\right)^2+\left(1-1\right)^2}\\
\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
AB=3\\[0.3cm]
BC=\sqrt{56}\equiv2\sqrt{14}\\[0.3cm]
AC=5\\
\end{array}\right."

ANSWER



"\boxed{AB=3\quad\quad BC=2\sqrt{14}\quad\quad AC=5}"

QUESTION 4.2

As in the previous problem, we calculate the lengths of the sides of the triangle with vertices "P\left(3,1,-1\right)" , "Q\left(2,0,3\right)" and "R\left(1,1,1\right)" , then check the fulfillment of the Pythagorean theorem.



"\left\{\begin{array}{l}
P\left(3,1,-1\right)\\
Q\left(2,0,3\right)\\
R\left(1,1,1\right)\\
\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
PQ=\sqrt{\left(3-2\right)^2+\left(1-0\right)^2+\left(3+1\right)^2}\\[0.3cm]
QR=\sqrt{\left(2-1\right)^2+\left(1-0\right)^2+\left(3-1\right)^2}\\[0.3cm]
PR=\sqrt{\left(3-1\right)^2+\left(1-1\right)^2+\left(1+1\right)^2}\\
\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
PQ=\sqrt{18}\equiv3\sqrt{2}\\[0.3cm]
QR=\sqrt{6}\\[0.3cm]
PR=\sqrt{8}\equiv2\sqrt{2}\\
\end{array}\right."

As "PQ>QR>PR" , only "PQ" can be the hypotenuse, then



"PQ^2\land QR^2+PR^2\\[0.3cm]
\left(\sqrt{18}\right)^2 \land \left(\sqrt{6}\right)^2+\left(\sqrt{8}\right)^2\\[0.3cm]
18\neq6+8"

Conclusion,



"\boxed{\Delta PQR-\text{not right triangle}}"

QUESTION 4.3

Let "\overrightarrow{u}=\left(0,1,1\right)" , "\overrightarrow{v}=\left(2,2,0\right)" and "\overrightarrow{w}=\left(−1,1,0\right)"

(a) As we know, if the angle between vectors "\overrightarrow{a}" and "\overrightarrow{b}" is "90^\circ", then "\overrightarrow{a}\cdot\overrightarrow{b}=0" .

In our case,



"\left\{\begin{array}{l}
\overrightarrow{u}=\left(0,1,1\right)\\[0.3cm]
 \overrightarrow{v}=\left(2,2,0\right)\\[0.3cm]
\overrightarrow{w}=\left(−1,1,0\right)\\
\end{array}\right.\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
\overrightarrow{u}\cdot\overrightarrow{v}=0\cdot2+1\cdot2+1\cdot0=2\neq0\\[0.3cm]
\overrightarrow{v}\cdot\overrightarrow{w}=-1\cdot2+1\cdot2+0\cdot0=0\\[0.3cm]
\overrightarrow{u}\cdot\overrightarrow{w}=-1\cdot0+1\cdot1+1\cdot0=1\neq0\\
\end{array}\right."

Conclusion,



"\boxed{\overrightarrow{v}\cdot\overrightarrow{w}=0\longrightarrow\overrightarrow{v}\perp\overrightarrow{w}}"

(b) As we know, the angle between the vectors can be found by the formula



"\cos\phi=\frac{\overrightarrow{u}\cdot\overrightarrow{w}}{\left\|\overrightarrow{u}\right\|\cdot\left\|\overrightarrow{w}\right\|}=\frac{1}{\sqrt{0^2+1^2+1^2}\cdot\sqrt{(-1)^2+1^2+0^2}}\\[0.3cm]
\cos\phi=\frac{1}{\sqrt{2}\cdot\sqrt{2}}=\frac{1}{2}\longrightarrow\boxed{\phi=60^\circ}"

QUESTION 4.5

Let "\overrightarrow{n}=(x,y,z)" is orthogonal to both "\overrightarrow{u}=(0,-1,-1)" and "\overrightarrow{v}=(1,0,-1)" , then

"\overrightarrow{n}\cdot\overrightarrow{u}=0" and "\overrightarrow{n}\cdot\overrightarrow{v}=0" .


"\left\{\begin{array}{l}
\overrightarrow{n}\cdot\overrightarrow{u}=0\cdot x+(-1)\cdot y+(-1)\cdot z=-y-z=0\\[0.3cm]
\overrightarrow{n}\cdot\overrightarrow{v}=1\cdot x+0\cdot y+(-1)\cdot z=x-z=0\\
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=-z\\x=z\\z=C
\end{array}\right.\longrightarrow\boxed{\overrightarrow{n}=\left(C,-C,C\right)}"

It remains to make the found vector unit, that is "\left\|\overrightarrow{n}\right\|=1" . In our case,



"\left\|\overrightarrow{n}\right\|=\sqrt{C^2+(-C)^2+C^2}=C\sqrt{3}=1\to\boxed{C=\frac{1}{\sqrt{3}}}"

Conclusion,



"\boxed{\overrightarrow{n}=\left(\frac{1}{\sqrt{3}},\frac{-1}{\sqrt{3}},\frac{1}{\sqrt{3}}\right)}"

QUESTION 4.4

Let "x < 0" . Find the vector "\overrightarrow{n}=\left(x,y,z\right)" that is orthogonal to all three vectors "\overrightarrow{u}=\left(1,1,-2\right)" , "\overrightarrow{v}=\left(−1,2,0\right)" and "\overrightarrow{w}=\left(−1,0,1\right)" . Then,



"\left\{\begin{array}{l}
\overrightarrow{n}\cdot\overrightarrow{u}=1\cdot x+1\cdot y+(-2)\cdot z=x+y-2z=0\\[0.3cm]
\overrightarrow{n}\cdot\overrightarrow{v}=-1\cdot x+2\cdot y+0\cdot z=-x+2y=0\\[0.3cm]
\overrightarrow{n}\cdot\overrightarrow{w}=-1\cdot x+0\cdot y+1\cdot z=-x+z=0\\
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
2y+y-2\cdot(2y)=0\rightarrow-y=0\\[0.3cm]
x=2y\\[0.3cm]
z=x=2y\\
\end{array}\right.\longrightarrow\\[0.3cm]
\boxed{\overrightarrow{n}=(0,0,0)}"